Explane why PCL5 is trigonal bipyramidal whereas IF5 is square pyramidal

Grade:11

1 Answers

Dwij
6 years ago
Okay I will tell you a method to find out the shapes. The Formula is 1/2(V+L+A-C) where V is valence electron of central atom, L is ligands(atoms attached to central atom, a is anion charge and c is cation charge)Using forumlaPCl5 = 1/2 ( 5+5+0+0) = 5Now if the number is 2 , it is sp hybridized, 3 sp2 4 sp3 5 sp3d 6 sp3d2 and so on. Here it is sp3d and structure of sp3d is trigonal bipyramidalIf5 = 1/2( 7+5+0+0)=6sp3d2 hybridized i.e square bipyramidal!